Operations Report – April 26th

Crew 315 Operations Report 26-04-2025

SOL: 6

Name of person filing report: Michael Andrews

Non-nominal systems: Suit Charger #7, Catalytic Heater, ScienceDome Radio Charger

Notes on non-nominal systems:

· (Unresolved from yesterday) It was noted that Suit #7 wasn’t charging, but further diagnosis from the Commander has isolated the issue to the #7 charger. I was able to observe that all chargers appear to work, so more formal diagnoses are required. (Note: this one has been on here a while, but we don’t any issues with chargers anymore, so this is low priority to work.)

· (Bonus project) The catalytic heater near the exercise bike in the habitat does not work. It’s an engineering project to see if this can get properly started during this mission.

· (Unresolved from yesterday) The ScienceDome radio chargers do not work. Mission support mentioned how this is a known issue. Radios will be swapped between the two structures, and I will disassemble the charger tomorrow.

ROVERS

Spirit rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Opportunity rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Curiosity rover used: Yes

Hours: 314.9 hours (1.1 hours used)

Beginning charge: 100%

Ending charge: 48%

Currently charging: Yes

Perseverance rover used: Yes

Hours: 329.9 hours (1.4 hours used)

Beginning charge: 100%

Ending charge: 56%

Currently charging: No

General notes on rovers: None

Summary of Hab operations: The team took this as a day of rest, performing only one EVA and working on other projects and personal time. The 5th (of 8) 100cameras project was held, and the team made cookies and spaghetti throughout the day. We also enjoyed Commander Laude’s show-and-tell of electronic hardware over the decades and an episode of Moonbase 8.

Water Use

Time of measurements: 1930
1) Per formula: 24.51 gallons
2) Smart Home Dashboard (see notes): 23.75 gallons

Water (static tank, remaining gallons): 303.3 gallons

Static tank pipe heater: Off

Static tank heater: Off

Toilet tank emptied: No

Summary of internet: Used to complete reports, learn languages, share team photos via Google drive, researching projects

Summary of suits and radios:

· EVA 11: Suits 8, 9, 10 and Radios 1, 4, 5

Summary of GreenHab operations: Watered plants and fungal project, kept the door open for most of the day to drive down temperature. Also harvested a small amount of basil for dinner.

WATER USE: 10.0 gallons

Heater: Yes

Supplemental light: 1900-2200

Harvest: 7g of basil

Summary of ScienceDome operations: I attempted to 3D scan the sample collected from EVA #10, but had difficulties with computer storage and had to abort the activity to be attempted later tonight.

Dual split (Heat or AC, On or Off): Off

Summary of RAM operations: No notable RAM activity today besides receiving a resupply of duct tape, flour, and raisin bran cereal. I also did want to report a broken plastic clamp that cracked when I was attempting to prototype the 3D scanning shroud earlier. I have tossed it in the RAM trash, and there are still 2 metal (and more resilient) clamps in inventory.

Summary of any observatory issues: None

Summary of health and safety issues: None

Questions, concerns and requests to Mission Support: None

Copyright © The Mars Society. All rights reserved. | Main Site